Log 10 x=1.5 then log 10 100g^5=????? plz solve this IIIII

Log10 x=1.5 then log10 100g^5=?????

plz solve this

 
 
IIIII

Grade:12th pass

4 Answers

Arun
5 years ago
Dear Hamid
 
What is g in this question, please make it clear. You can also attach an image.
I will be happy to help you.
Hamid Ali Shah
5 years ago
it is x sorry my mistake.
Arun
5 years ago
Dear student
 
log x = 1.5
x = 10^1.5 = sqrt(1000)
Now 
log 100 * sqrt(1000) = log 100 + log sqrt(1000) = 2 + 3/2 = 3.5
Regards
Arun
Aryan
one year ago
log x = 1.5
x = 10^1.5 = sqrt(1000)
Now 
log 100 * sqrt(1000) = log 100 + log sqrt(1000) = 2 + 3/2 = 3.5